मराठी

Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.

Advertisements
Advertisements

प्रश्न

Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.

पर्याय

  • `186/190`

  • `187/190`

  • `188/190`

  • `18/(""^20C_3)`

MCQ
रिकाम्या जागा भरा
Advertisements

उत्तर

Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive `187/190`.

Explanation:

Set of three consecutive numbers from 1 to 20 are 1, 2, 3; 2, 3, 4; 3, 4, 5; ..., 18, 19, 20.

So, the probability that the numbers are consecutive

= `18/(""^20C_3)`

= `18/((201)/(3!17!))`

= `(18*3!*17!)/(20!)`

= `(18 xx 3 xx 2 xx 17!)/(20 xx 19 xx 18 xx 17!)`

= `(3 xx 2)/(20 xx 19)`

= `3/190`

∴ P(three numbers are not consecutive) = `1 - 3/190 = 187/190`

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 16: Probability - Exercise [पृष्ठ २९९]

APPEARS IN

एनसीईआरटी एक्झांप्लर Mathematics Exemplar [English] Class 11
पाठ 16 Probability
Exercise | Q 19 | पृष्ठ २९९

व्हिडिओ ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्‍न

Which of the following can not be valid assignment of probabilities for outcomes of sample space S = {ω1, ω2,ω3,ω4,ω5,ω6,ω7}

Assignment ω1 ω2 ω3 ω4 ω5 ω6 ω7
(a) 0.1 0.01 0.05 0.03 0.01 0.2 0.6
(b) `1/7` `1/7` `1/7` `1/7` `1/7` `1/7` `1/7`
(c) 0.1 0.2 0.3 0.4 0.5 0.6 0.7
(d) –0.1 0.2 0.3 0.4 -0.2 0.1 0.3
(e) `1/14` `2/14` `3/14` `4/14` `5/14` `6/14` `15/14`

A card is selected from a pack of 52 cards.

  1. How many points are there in the sample space?
  2. Calculate the probability that the card is an ace of spades.
  3. Calculate the probability that the card is
    1. an ace
    2. black card.

A fair coin with 1 marked on one face and 6 on the other and a fair die are both tossed. Find the probability that the sum of numbers that turn up is (i) 3 (ii) 12


Three coins are tossed once. Find the probability of getting

  1. 3 heads
  2. 2 heads
  3. at least 2 heads
  4. at most 2 heads
  5. no head
  6. 3 tails
  7. exactly two tails
  8. no tail
  9. atmost two tails.

Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6


A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that all will be blue?


A and B throw a pair of dice. If A throws 9, find B's chance of throwing a higher number.

 

Two unbiased dice are thrown. Find the probability that the total of the numbers on the dice is greater than 10.

 

Two unbiased dice are thrown. Find the probability that  neither a doublet nor a total of 8 will appear


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that all the three balls are blue balls 


If a letter is chosen at random from the English alphabet, find the probability that the letter is  a vowel .


In a lottery, a person chooses six different numbers at random from 1 to 20, and if these six numbers match with six number already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game?


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iv)
\[\frac{1}{14}\]
\[\frac{2}{14}\]
\[\frac{3}{14}\]
\[\frac{4}{14}\]
\[\frac{5}{14}\]
\[\frac{6}{14}\]
\[\frac{15}{14}\]

In a single throw of three dice, find the probability of getting the same number on all the three dice.


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that: all 10 are defective


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that none is defective


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 1, 2, 3, 4 or 5


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 5, 15, 25, or 35


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that C will be selected?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine `P(B ∩ barC)`


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine `P(A ∩ barB)`


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(B ∩ C)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine Probability of exactly one of the three occurs


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are white


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are red


If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together


If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together


In a non-leap year, the probability of having 53 tuesdays or 53 wednesdays is ______.


A single letter is selected at random from the word ‘PROBABILITY’. The probability that it is a vowel is ______.


If A and B are two candidates seeking admission in an engineering College. The probability that A is selected is .5 and the probability that both A and B are selected is at most .3. Is it possible that the probability of B getting selected is 0.7?


The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×